### v2

# Multi-Partite Output Regulation of Multi-Agent Systems

Abstract
This article proposes a simple, graph-independent perspective on partitioning the node set of a graph and provides multi-agent systems (MASs) with objectives beyond cooperation and bipartition. Specifically, we first introduce the notion of $k$-partition transformation to achieve any desired partition of the nodes. Then, we use this notion to formulate the multi-partite output regulation problem (MORP) of heterogeneous linear MASs, which comprises the existing cooperative output regulation problem (CORP) and bipartite output regulation problem (BORP) as subcases. The goal of the MORP is to design a distributed control law such that each follower that belongs to the same set in the partition asymptotically tracks a scalar multiple of the reference while ensuring the internal stability of the closed-loop system. It is shown that the necessary and sufficient conditions for the solvability of the MORP with a feedforward-based distributed control law follow from the CORP and lead to the first design strategy for the control parameters. However, it has a drawback in terms of scalability due to a partition-dependent condition. We prove that this condition is implied by its partition-independent version under a mild structural condition. This implication yields the second design strategy that is much more scalable than the first one. Finally, an experiment is conducted to demonstrate the MORP’s flexibility, and two numerical examples are provided to illustrate its generality and compare both design strategies regarding scalability.
Index Terms:
Cooperative control, distributed control, output regulation, linear matrix equations, multi-agent system.
I Introduction
I-A Motivation and Literature Review
Considering the studies on distributed control of multi-agent systems (MASs), two main frameworks become distinguishable: cooperative [1] and bipartite [2],[3]. The most recognized problems of both frameworks include consensus, formation tracking, and output regulation. Regardless of the problem, MASs have a common objective in the cooperative framework, whereas they potentially have two opposed objectives in the bipartite framework.
This article is mainly motivated by the need for a flexible framework that extends beyond cooperation and bipartition within MASs to accommodate multiple, shifting mission objectives in adverse operating environments. For example, consider networked uninhabited aerial vehicles (UAVs) tasked with suppressing and destroying enemy air defense [4]. During this operation, having the flexibility to achieve arbitrarily changing formations is paramount to avoid radars while maximizing damage. Given that the output regulation problems for MASs allow high-order nonidentical agent dynamics and contain typical cooperative control problems such as leader-following consensus and formation tracking, we propose the multi-partite output regulation problem (MORP) to enhance tactical flexibility of MASs. The MORP includes the cooperative output regulation problem (CORP) and the bipartite output regulation problem (BORP) as special cases.
Analogous to the output regulation problem, the CORP has been mainly treated using feedforward[5, 6, 7, 8, 9] and internal model [10, 11, 12, 13, 14] approaches. In the former, the feedforward gain of each agent is based on the solution of the regulator equations, which are linear matrix equations (LMEs) determined by the exosystem and agent dynamics, making it not robust to parameter uncertainties. While the latter is known to be robust against small parameter variations, it cannot be applied when the transmission zeros condition does not hold.
Compared to the CORP, fewer studies exist on the BORP. It is tackled using the feedforward approach in [15, 16, 17] and the internal model approach in [18]. In the CORP (BORP), the feedforward approach, with controller state exchange of neighboring agents, uses a distributed observer to provide the estimated state (estimated state or its additive inverse) of the exosystem to every agent.
Apart from the bipartite framework, there have been efforts to increase the number of objectives in MASs. The notable ones are cluster consensus [19, 20, 21], scaled consensus [22],[23] and kernel manipulation of the Laplacian matrix [24],[25]. Yet, those studies are limited to consensus problems over first or second order agent dynamics. The cluster problem is also extended to high-order heterogeneous MASs [26, 27]. The proposed MORP differs from the cluster consensus in two main aspects: First, the partitioning in the MORP is independent of the underlying graph, allowing each agent to determine its set in the partition, and hence, its objective, whereas the clustering in cluster consensus is graph-dependent, preventing each agent from self-determining its cluster. Second, while the MORP requires only one leader to generate multiple objectives, the cluster consensus requires at least two leaders to yield more than one objective.
I-B Contribution
This article formulates a general distributed control problem for heterogeneous MASs. To solve this problem, it provides two design strategies with comparable advantages and disadvantages for a feedforward-based distributed control law involving a distributed observer.
The partition of the node set of a graph in the bipartite framework is graph-dependent (e.g., see Remark 2.1 and Equation (8) in [18]). In particular, each agent needs to know the graph’s adjacency matrix to determine the set in which it lies in the bipartition, and hence, track the reference or its additive inverse. To render the partition of the node set independent of the graph, we introduce the notion of $k$-partition transformation. This not only allows the node set to admit up to $N$-partition, where $N$ is the cardinality of the node set, but also provides each agent with the flexibility to self-determine the set in which it lies in the $k$-partition, and hence, which scalar multiple of the reference to track.
By leveraging this notion, we formulate the MORP, which includes the CORP and BORP as special cases. To this end, we consider a heterogeneous MAS in the form
$\dot{x}_{i}$ $=A_{i}x_{i}+B_{i}u_{i}+E_{i}v$
$y_{i}$ $=C_{i}x_{i}+D_{i}u_{i}+G_{i}v,\quad i=1,\dots,N$ (1)
where $x_{i}\in\mathbb{R}^{n_{i}}$ is the state, $u_{i}\in\mathbb{R}^{m_{i}}$ is the control input, and $y_{i}\in\mathbb{R}^{p_{i}}$ is the output of subsystem $i$. Also, $v\in\mathbb{R}^{n_{0}}$ is the exogenous signal generated by the following exosystem
$\dot{v}=A_{0}v.$ (2)
This autonomous linear system yields the disturbances $E_{i}v$ and $G_{i}v$ and the reference denoted by $-F_{i}v$ for subsystem $i$. The goal of the MORP is to design a distributed control law such that each subsystem in the same set within the partition asymptotically tracks a common scalar multiple of the reference (i.e., a $k$-partition transformation term multiple of the reference) and rejects the disturbance while ensuring the internal stability of the closed-loop system.
Although the MORP can be cast as the CORP, and hence, the necessary and sufficient solvability conditions of the MORP follow from the CORP, the regulator equations become partition-dependent (see Condition 3). Consequently, the immediate design strategy for control parameters, which is called the first design strategy, requires each subsystem to recompute a solution pair for the regulator equations each time the $k$-partition transformation term is updated. This is a drawback that does not exist in the CORP.
An intriguing question arises from the drawback above: Is it possible to have a design strategy that eliminates the recomputation of a solution for the regulator equations whenever the $k$-partition transformation term changes? Theorem 2 paves the way for an affirmative answer under a mild structural condition. Specifically, it shows that the solvability of the partition-independent regulator equations (3) and the introduced LME (10), which solely depends on the subsystem data, ensures the existence of a solution pair to the partition-dependent regulator equations (3). More importantly, Theorem 2 provides an analytical formula (2) for such a pair in terms of $s_{i}$ and the solutions of (3) and (10). This leads to the second design strategy that does not require recomputing a solution to any equations when the $k$-partition transformation term changes. Accordingly, it is significantly more scalable than the first one. The discussion on the proposed design strategies’ scalability and conservatism is summarized in Table I by referring to the corresponding conditions and results of the paper. An experiment is conducted to demonstrate the MORP’s flexibility in accommodating shifting mission objectives. Two numerical examples are also provided to showcase its generality and compare both design strategies regarding scalability.
I-C Notation
The real part of a complex number $\lambda$ is denoted by $\mathrm{Re}(\lambda)$. The closed right (left) half complex plane is denoted by $\mathrm{CRHP}$ ($\mathrm{CLHP}$). The open right half complex plane is denoted by $\mathrm{ORHP}$. We write $I_{n}$ for the $n\times n$ identity matrix, $0_{n\times m}$ or 00 for the $n\times m$ zero matrix, $\mathrm{diag}(w_{1},\ldots,w_{n})$ for the diagonal matrix with scalar entries $w_{1},\ldots,w_{n}$ on its diagonal, and $\otimes$ for the Kronecker product. The spectrum of a square matrix $X\in\mathbb{R}^{n\times n}$ is denoted by $\mathrm{spec}(X)$. The matrix obtained by replacing each entry of $X$ with its absolute value is denoted by $|X|$. The image of a matrix $Z\in\mathbb{R}^{n\times m}$ is denoted by $\mathrm{im}\hskip 0.85355ptZ$.
A (weighted) signed directed graph ${\mathcal{G}}$ is a triple ${\mathcal{G}}=(\mathcal{N},\mathcal{E},\mathcal{A})$, where $\mathcal{N}=\left\{1,\ldots,N\right\}$ is the node set, $\mathcal{E}\subseteq\mathcal{N}\times\mathcal{N}$ is the edge set, and $\mathcal{A}=[a_{ij}]\in\mathbb{R}^{N\times N}$ is the adjacency matrix whose entries are determined by the rule: for any $j,i\in\mathcal{N}$, $a_{ij}\neq 0$ if, and only if, $(j,i)\in\mathcal{E}$. Graphs with self-loops are not considered in this article; that is, $a_{ii}=0$ for $i=1,\ldots,N$. A graph is completely specified by its adjacency matrix $\mathcal{A}$, hence, the graph corresponding to $\mathcal{A}$ is denoted as $\mathcal{G}(\mathcal{A})$. A signed directed graph ${\mathcal{G}(\mathcal{A})}$ is called an unsigned directed graph if $a_{ij}\geq 0$ for $i,j=1,\ldots,N$. The in-degree and Laplacian matrices of a signed directed graph $\mathcal{G}(\mathcal{A})$, denoted by $\mathcal{D}$ and $\mathcal{L}$, are defined as $\mathcal{D}=\mathrm{diag}(d_{1},\ldots,d_{N})$ with $d_{i}=\sum_{j=1}^{N}|a_{ij}|$ for $i=1,\ldots,N$ and $\mathcal{L}=\mathcal{D}-\mathcal{A}$.
II Problem Formulation
This section introduces the notion of $k$-partition transformation to achieve any desired partition for the node set of a given graph. It then uses this notion to formulate the MORP of heterogeneous linear MASs.
II-A Arbitrary Partition of Nodes
We first recall the $k$-partition of sets from combinatorics.
Definition 1 (Sections 1.10, 5.1, and 5.4 in [28])
Let $T$ be a nonempty finite set of $N$ elements. A collection $\mathcal{T}$ of $1\leq k\leq N$ nonempty subsets of $T$ is called a $k$-partition of $T$ if all the sets in $\mathcal{T}$ are mutually disjoint and if their union equals $T$. The number of $k$-partitions of $T$ is called the Stirling number of the second kind. The number of all partitions of $T$ is called the Bell number.
Then, we introduce the $k$-partition transformation concept. This generalizes the gauge transformation (i.e., the signature matrix) used in the bipartite framework [2, 29, 18].
Definition 2
A matrix $S=\mathrm{diag}(s_{1},\ldots,s_{N})\in\mathbb{R}^{N\times N}$ with exactly $1\leq k\leq N$ distinct entries on the diagonal is called a $k$-partition transformation, where each $s_{i}$ is called a $k$-partition transformation term. A 1-partition or 2-partition transformation $S$ is a gauge transformation if $s_{i}\in\{-1,1\}$ for $i=1,\ldots,N$.
Given a signed directed graph, the bipartite framework partitions the nodes according to the signs of the adjacency matrix entries (i.e., the edge weights). This yields at most $2$-partition of nodes. In contrast, we partition the nodes according to a $k$-partition transformation. Since a $k$-partition transformation can be chosen for any $k\in\{1,\ldots,N\}$, the nodes can admit not only 1-partition or 2-partition but also more than 2-partition. In fact, there are infinitely many $k$-partition transformations so that the nodes can be partitioned in the Bell number of ways because $k$-partition transformation terms can be any real number. We now formally discuss the $k$-partition of nodes.
Let $\mathcal{G}(\mathcal{A})$ be a signed directed graph, and let $S$ be a $k$-partition transformation. Then, there exist $k$ positive integers $i_{1},\ldots,i_{k}$ such that $s_{i_{1}},\ldots,s_{i_{k}}$ are $k$ distinct $k$-partition transformation terms. For each $p\in\left\{1,\ldots,k\right\}$, let
$\mathcal{N}_{p}=\left\{j\in\mathcal{N}~{}|~{}s_{j}=s_{i_{p}}\right\}.$ (3)
Define the collection $\mathcal{C}=\left\{\mathcal{N}_{1},\ldots,\mathcal{N}_{k}\right\}$. Lemma 1 verifies that it is a $k$-partition of the node set $\mathcal{N}$.
Lemma 1
The collection $\mathcal{C}$ has the following properties:
  1. (i)
    If $\mathcal{N}_{p}\in\mathcal{C}$, then $\mathcal{N}_{p}\neq\emptyset$.
  2. (ii)
    If $\mathcal{N}_{p}\in\mathcal{C}$ and $\mathcal{N}_{r}\in\mathcal{C}$ with $p\neq r$, then $\mathcal{N}_{p}\cap\mathcal{N}_{r}=\emptyset$.
  3. (iii)
    $\bigcup_{\mathcal{N}_{p}\in\mathcal{C}}\mathcal{N}_{p}=\mathcal{N}$.
Proof:
(i) Clearly, $i_{p}\in\mathcal{N}_{p}$. (ii) Let $\mathcal{N}_{p}\in\mathcal{C}$ and $\mathcal{N}_{r}\in\mathcal{C}$ with $p\neq r$, but assume for contradiction that $\mathcal{N}_{p}\cap\mathcal{N}_{r}\neq\emptyset$. Then, let $j\in\mathcal{N}_{p}\cap\mathcal{N}_{r}$; hence, $j\in\mathcal{N}_{p}$ and $j\in\mathcal{N}_{r}$. By definition, $s_{j}=s_{i_{p}}$ and $s_{j}=s_{i_{r}}$. Thus, $s_{i_{p}}=s_{i_{r}}$, which contradicts $s_{i_{p}}$ and $s_{i_{r}}$ being distinct. (iii) Since every $\mathcal{N}_{p}\in\mathcal{C}$ is a subset of $\mathcal{N}$, the inclusion $\bigcup_{\mathcal{N}_{p}\in\mathcal{C}}\mathcal{N}_{p}\subseteq\mathcal{N}$ holds. Let $j\in\mathcal{N}$. Then, there exists a $p\in\left\{1,\ldots,k\right\}$ such that $s_{j}=s_{i_{p}}$ due to the fact that $S$ has exactly $k$ distinct diagonal entries. Hence, $j\in\mathcal{N}_{p}$ for some $\mathcal{N}_{p}\in\mathcal{C}$. This proves that the inclusion $\mathcal{N}\subseteq\bigcup_{\mathcal{N}_{p}\in\mathcal{C}}\mathcal{N}_{p}$ holds. ∎
Remark 1
Owing to Definition 2 and Lemma 1, the $k$-partition transformations can achieve arbitrary partition of nodes. For example, let $\mathcal{G}(\mathcal{A})$ be a graph with $5$ nodes. For $k=1,2,3,4,5$, there are infinitely many $k$-partition transformations that can obtain $1,15,25,10,1$ number of $k$-partitions of the nodes, respectively. These numbers correspond to the Stirling numbers of the second kind for respective values of $k$ (e.g., see Section 5.1 in [28]). Their sum is $52$, which is the associated Bell number.
II-B MORP
As in the context of the CORP and BORP, the subsystems of (I-B), considered followers, and the exosystem (2), considered the leader, constitute a MAS of $N+1$ agents. To model the information exchange between $N$ followers, we use a time-invariant signed directed graph ${\mathcal{G}}(\mathcal{A})$ without self-loops as $\mathcal{N}=\left\{1,\ldots,N\right\}$, where node $i\in\mathcal{N}$ corresponds to follower $i$, and for each $j,i\in\mathcal{N}$, we put $(j,i)\in\mathcal{E}$ if, and only if, follower $i$ has access to the information of follower $j$. The leader is included in the information exchange model by augmenting the graph ${\mathcal{G}}(\mathcal{A})$. Specifically, let $\mathcal{G}(\bar{\mathcal{A}})$ be an augmented signed direct graph with $\bar{\mathcal{N}}=\mathcal{N}\cup\{0\}$, $\bar{\mathcal{E}}=\mathcal{E}\cup\mathcal{E}^{\prime}$, where $\mathcal{E}^{\prime}\subseteq\{(0,i)\ |\ i\in\mathcal{N}\}$, and $\bar{\mathcal{A}}\in\mathbb{R}^{(N+1)\times(N+1)}$. Here, node 00 corresponds to the leader and for any $i\in\mathcal{N}$, we put $(0,i)\in\mathcal{E}^{\prime}$ if, and only if, follower $i$ has access to the information of the leader. For any $i\in\mathcal{N}$, the pinning gain $f_{i}>0$ if $(0,i)\in\mathcal{E}^{\prime}$ and $f_{i}=0$ otherwise. The pinning gain matrix is defined by $\mathcal{F}=\mathrm{diag}(f_{1},\ldots,f_{N})$.
A control law that relies on the information exchange modeled by an augmented signed directed graph $\mathcal{G}(\bar{\mathcal{A}})$ is called a distributed control law. The closed-loop system consists of (I-B) and the distributed controller. We now formulate the MORP.
Problem 1 (MORP)
Given the heterogeneous MAS composed of (I-B) and (2), an augmented signed directed graph $\mathcal{G}(\bar{\mathcal{A}})$, and a $k$-partition transformation $S$, find a distributed control law such that
  1. (i)
    The closed-loop system matrix is Hurwitz;
  2. (ii)
    For any initial state of the exosystem and closed-loop system, the tracking error of each $i\in\mathcal{N}$ defined by
    $e_{i}=y_{i}+s_{i}F_{i}v$ (4)
    satisfies $\lim_{t\to\infty}e_{i}(t)=0$.
Remark 2
The MORP includes the CORP and BORP objectives: If $S=I_{N}$, the MORP reduces to the CORP (e.g., see Definition 1 in [5]). If $S$ is a gauge transformation, the MORP reduces to the BORP (e.g., see Problem 2.1 in [18]).
III Solvability of the MORP
This section first observes that the MORP can be realized as the CORP. We then consider one of the feedforward-based distributed control laws solving the CORP. Based on this control law and the observation, necessary and sufficient conditions for the solvability of the MORP follow from the CORP, resulting in the first design strategy for control parameters. Though this strategy is straightforward, it has a drawback due to the $k$-partition transformation dependence of the regulator equations. The section concludes with a discussion of this drawback.
III-A Distributed Control Law
[…]
$\dot{x}_{i}$ $=A_{i}x_{i}+B_{i}u_{i}+E_{i}v$
$e_{i}$ $=C_{i}x_{i}+D_{i}u_{i}+\tilde{F}_{i}v,\quad i=1,\ldots,N.$ (5)
Then, the following result is immediate.
Lemma 2
Any distributed control law solving the CORP for the MAS composed of (III-A) and (2) over an augmented unsigned directed graph solves the MORP over the same graph.
In the light of Lemma 2, we consider the following feedforward-based distributed control law [^1] With the distributed measurement output feedback control law in Equation (8.14) of [1], one can arrive at a result similar to Theorem 1 (ii) under the additional detectability assumptions. The resulting design strategy will have the drawback in Remark 7. In this case, Theorem 2 is still a remedy. proposed in [5]
$\dot{\eta}_{i}$ $=$ $A_{0}\eta_{i}+\mu\Big{(}\sum_{j=1}^{N}|a_{ij}|(\eta_{j}-\eta_{i})+f_{i}(v-\eta_{i})\Big{)}$
$u_{i}$ $=$ $K_{1i}x_{i}+K_{2i}\eta_{i},\quad i=1,\ldots,N$ (6)
where $\eta_{i}\in\mathbb{R}^{n_{0}}$ is the estimate of $v$. The state equation in (III-A) is called a distributed observer. Moreover, $\mu\in\mathbb{R}$, $K_{1i}\in\mathbb{R}^{m_{i}\times n_{i}}$, and $K_{2i}\in\mathbb{R}^{m_{i}\times n_{0}}$ are control parameters to be designed. Compared to the distributed observer in [5], the one in (III-A) uses the absolute value of the adjacency matrix entries to be readily applicable even if the given graph is signed.
III-B MORP: Necessary and Sufficient Conditions
The following conditions will be referred to for the solvability of Problem 1.
Condition 1
The inclusion $\mathrm{spec}(A_{0})\subsetneq\mathrm{CRHP}$ holds.
Condition 2
For any $i\in\mathcal{N}$, the pair $(A_{i},B_{i})$ is stabilizable.
Condition 3
For any $i\in\mathcal{N}$, there exists a pair $(X_{i},U_{i})$ that satisfies the regulator equations
$X_{i}A_{0}$ $=$ $A_{i}X_{i}+B_{i}U_{i}+E_{i}$
00 0 $=$ $C_{i}X_{i}+D_{i}U_{i}+G_{i}+s_{i}F_{i}.$ (7)
Condition 4
The augmented signed directed graph $\mathcal{G}(\bar{\mathcal{A}})$ contains a directed spanning tree [^2] By the definition of the augmented signed directed graph given in Section II-B, the root of any directed spanning tree in $\mathcal{G}(\bar{\mathcal{A}})$ is necessarily node 00..
Remark 3
Conditions 1, 2, and 4 are standard while tackling both the CORP and BORP (e.g., see [5, 17]). The BORP literature imposes the structural balance condition on either $\mathcal{G}(\bar{\mathcal{A}})$ (e.g., see Assumption 3.1 in [15]) or $\mathcal{G}(\mathcal{A})$ (e.g., see Assumption 5 in [17]). Yet, this article does not require such a condition because the distributed observer in (III-A) uses the absolute value of the adjacency matrix entries.
Remark 4
The $k$-partition transformation term $s_{i}$ is incorporated into (3) due to Lemma 2. Except for this term, Condition 3 is standard for the studies investigating the CORP with the feedforward approach (e.g., see [5, 8]). The articles [17, 15] studying the BORP with the feedforward approach consider the disturbance-free (i.e., $E_{i}=0$ and $G_{i}=0$) and direct feedthrough-free (i.e., $D_{i}=0$) follower dynamics. However, introducing a gauge transformation term in the regulator equations for each follower can extend their distributed controllers to take the effect of disturbances and direct feedthrough into account. To avoid such modification, the authors in [16] assume that agents in different sets in partition are subject to disturbances that are additive inverses of each other. This can be impractical in real-world applications, for instance, consider networked UAVs operating in the same environment. Thus, this article does not make that assumption about disturbances.
We now provide the necessary and sufficient conditions for the solvability of the MORP.
Theorem 1
The following statements are true:
  1. (i)
    Suppose Condition 1 holds. If Problem 1 is solvable by the distributed control law (III-A), then Conditions 24 hold.
  2. (ii)
    Under Conditions 24, Problem 1 is solvable by the distributed control law (III-A).
Proof:
The proof can be conducted by following the procedure in the proof of Theorem 1 in [30]. ∎
Remark 5
Conditions 24 are sufficient for the solvability of the MORP by the distributed control law (III-A). Under Condition 1, they are also necessary.
III-C Discussion: Synthesis of Control Parameters
This subsection presents the first design strategy for the control parameters and highlights an associated drawback. Let $\mathcal{L}_{u}$ denote the Laplacian matrix of the unsigned directed graph $\mathcal{G(|\mathcal{A}|)}$ and $\mathcal{H}=\mathcal{L}_{u}+\mathcal{F}$. Let $\lambda_{i}(\mathcal{H})$ and $\lambda_{j}(A_{0})$ denote the eigenvalues of $\mathcal{H}$ and $A_{0}$ for $i=1,\ldots,N$, $j=1,\ldots,n_{0}$.
Remark 6
Let Conditions 24 hold. The constructive procedure in the proof of Theorem 1 (ii) yields the following design steps for the parameters $\mu$, $K_{1i}$, and $K_{2i}$:
  1. (i)
    Select $\mu$ according to the inequality333 The sufficient and necessary conditions for the Hurwitzness of the matrix $A_{\mu}=(I_{N}\otimes A_{0})-\mu(\mathcal{H}\otimes I_{n_{0}})$, which is the system matrix of the distributed observer in compact form, are used in the proof of Theorem 1. These conditions are given in Lemma 4 of [30]. However, the lower bound on $\mu$ in (8), due to Lemma 3 given in Appendix, has no conservatism compared to the bound in Lemma 4 of [30]. Another lower bound is also provided in Lemma 3.2 of [1]. But, one can easily construct a counterexample to the first statement of that lemma by considering a Hurwitz $A_{0}$.
    $\mu>\max_{\begin{subarray}{c}j\in\{1,\ldots,n_{0}\}\\ i\in\{1,\ldots,N\}\end{subarray}}\frac{\mathrm{Re}(\lambda_{j}(A_{0}))}{\mathrm{Re}(\lambda_{i}(\mathcal{H}))}.$ (8)
  2. (ii)
    For each $i\in\mathcal{N}$, design $K_{1i}$ such that $A_{i}+B_{i}K_{1i}$ is Hurwitz.
  3. (iii)
    For each $i\in\mathcal{N}$, find a pair $(X_{i},U_{i})$ that satisfies the regulator equations (3).
  4. (iv)
    For each $i\in\mathcal{N}$, let $K_{2i}=U_{i}-K_{1i}X_{i}$.
As $\mathrm{spec}(\mathcal{H})\subsetneq\mathrm{ORHP}$ under Condition 4 (e.g., see Lemma 1 in [5]), any positive $\mu$ satisfies the inequality (8) if $\mathrm{spec}(A_{0})\subsetneq\mathrm{CLHP}$. Thus, the design of $\mu$ is independent of the eigenvalues of $\mathcal{H}$ (i.e., the spectral property of $\mathcal{G}(\bar{\mathcal{A}})$) when the exosystem generates a linear combination of constant signals, sinusoidal signals, polynomial signals, for instance, ramp signals, and exponentially converging signals. These signals cover references and disturbances encountered in most multi-agent systems (e.g., see the applications in [3, 6, 9, 31]).
Remark 7
Though the design based on Remark 6 is straightforward to apply, it has a drawback in that the pair $(X_{i},U_{i})$ for each follower depends on $s_{i}$. Therefore, whenever the $k$-partition transformation is updated, it necessitates each follower to recompute a solution pair for the regulator equations (3) unless its $s_{i}$ remains unchanged. Hence, given a finite set of $k$-partition transformation terms with $M$ elements, each follower computes $M$ solution pairs to (3). This drawback can render the design strategy impractical in some applications. For example, consider low-cost networked UAVs trying to avoid radars. In this scenario, the set of $k$-partition transformation terms for each follower may not be known before the operation. Thus, it may be desirable in terms of computational cost to find a solution pair for the regulator equations (3) once and use it throughout the operation.
The following section, motivated by the discussion in Remark 7, seeks an answer to the question posed in Section I-B.
IV MORP: Partition-Independent Solvability
This section first provides the partition-independent sufficient conditions for the solvability of the MORP. For each $i\in\mathcal{N}$, these conditions include the partition-independent regulator equations, as in the CORP with the feedforward approach, and an LME depending only on $B_{i}$, $E_{i}$, $D_{i}$, and $G_{i}$. The constructive nature of the result yields the second design strategy that eliminates the drawback discussed in Remark 7. Lastly, we reveal that the intoduced LME’s solvability is equivalent to an easily testable mild structural condition.
IV-A MORP: Partition-Independent Sufficient Conditions
We modify Condition 3 by removing the $k$-partition transformation term in (3).
Condition 3
For any $i\in\mathcal{N}$, there exists a pair $(X_{i},U_{i})$ that satisfies the partition-independent regulator equations
$X_{i}A_{0}$ $=$ $A_{i}X_{i}+B_{i}U_{i}+E_{i}$
00 0 $=$ $C_{i}X_{i}+D_{i}U_{i}+G_{i}+F_{i}.$ (9)
Theorem 2 not only shows that the solvability of the partition-independent regulator equations (3) and the LME (10) ensures the solvability of the partition-dependent regulator equations (3) but also provides a solution pair.
Theorem 2
If Condition 3 holds and if, for any $i\in\mathcal{N}$, there exists a solution to the following LME
$\begin{bmatrix}B_{i}\\ D_{i}\end{bmatrix}Y_{i}=\begin{bmatrix}E_{i}\\ G_{i}\end{bmatrix}$ (10)
then Condition 3 holds. In particular, for any $i\in\mathcal{N}$, the pair $(\tilde{X}_{i},\tilde{U}_{i})$ given by
$\tilde{X}_{i}$ $=$ $s_{i}X_{i}$
$\tilde{U}_{i}$ $=$ $s_{i}(U_{i}+Y_{i})-Y_{i}$ (11)
satisfies the partition-dependent regulator equations (3).
Proof:
Fix $i\in\mathcal{N}$. Let $(X_{i},U_{i})$ be a pair that satisfies the partition-independent regulator equations (3). Also, let $Y_{i}$ be a solution to the LME (10). By the pair $(\tilde{X}_{i},\tilde{U}_{i})$ in (2),
$A_{i}\tilde{X}_{i}+B_{i}\tilde{U}_{i}+E_{i}$ $=s_{i}(A_{i}X_{i}+B_{i}U_{i}+B_{i}Y_{i})-B_{i}Y_{i}+E_{i}$
$=s_{i}(A_{i}X_{i}+B_{i}U_{i}+E_{i})$
$=s_{i}X_{i}A_{0}=\tilde{X}_{i}A_{0}$ (12)
where the second equation follows from $B_{i}Y_{i}=E_{i}$, the third equation is due to the fact that $A_{i}X_{i}+B_{i}U_{i}+E_{i}=X_{i}A_{0}$, and the fourth equation is a consequence of $\tilde{X}_{i}=s_{i}X_{i}$. By the pair $(\tilde{X}_{i},\tilde{U}_{i})$ in (2), we further have
$\hskip 2.84544ptC_{i}\tilde{X}_{i}+D_{i}\tilde{U}_{i}+G_{i}+s_{i}F_{i}$
$=$ $\hskip 2.84544pts_{i}\left(C_{i}X_{i}+D_{i}(U_{i}+Y_{i})+F_{i}\right)-D_{i}Y_{i}+G_{i}$
$=$ $\hskip 2.84544pts_{i}(C_{i}X_{i}+D_{i}U_{i}+G_{i}+F_{i})=0$ (13)
where the second equation follows from $D_{i}Y_{i}=G_{i}$ and the third equation is a result of $C_{i}X_{i}+D_{i}U_{i}+G_{i}+F_{i}=0$. We now conclude from (IV-A) and (IV-A) that the pair $(\tilde{X}_{i},\tilde{U}_{i})$ in (2) satisfies the partition-dependent regulator equations (3). Hence, the proof is over. ∎
Remark 8
The converse of the first statement in Theorem 2 is not true. To show this, consider a MAS consisting of one leader and one follower with system parameters $A_{0}=G_{1}=0$, $A_{1}=B_{1}=C_{1}=D_{1}=1$, $E_{1}=-2$, $F_{1}=-1$, and $k$-partition transformation term $s_{1}=2$. It can be verified that the pair $(1,1)$ satisfies the partition-dependent regulator equations (3). We assume for contradiction that there is a pair $(X_{1},U_{1})$ that satisfies the partition-independent regulator equations (3) and a solution $Y_{1}$ to the LME (10). Then, (3) yields $X_{1}+U_{1}=2$ and $X_{1}+U_{1}=1$. Hence, $1=2$, a contradiction. We have just proved that the converse of the first statement in Theorem 2 is not true. One can also show that the LME (10) does not have a solution for the considered example.
In conjunction with Theorem 1 (ii), Theorem 2 leads to the following partition-independent sufficient conditions for the solvability of the MORP.
Corollary 1
Under Conditions 2, 3, and 4, Problem 1 is solvable by the distributed control law (III-A) if, for any $i\in\mathcal{N}$, there is a solution $Y_{i}$ to the LME (10).
The rest of this subsection recalls a well-known alternative sufficient condition for Condition 3, highlights the importance of Theorem 2 in design, and compare the sufficient conditions.
Remark 9
Alongside Theorem 2, we know from Theorem 1.9 in [32] that another partition-independent sufficient condition for Condition 3 is that, for any $i\in\mathcal{N}$, the rank condition [^4] It is known as the transmission zeros condition (e.g., see Remark 1.11 in [32]) if the pair $(A_{i},B_{i})$ is controllable and the pair $(A_{i},C_{i})$ is observable.
$\mathrm{rank}\begin{bmatrix}A_{i}-\lambda_{j}(A_{0})&B_{i}\\ C_{i}&D_{i}\end{bmatrix}=n_{i}+p_{i},\quad j=1,\ldots,n_{0}$ (14)
holds. Similar to the conditions in Theorem 2, this sufficient condition is not necessary for Condition 3 to hold. To see this, consider the example in Remark 8. What significantly distinguishes Theorem 2 from this sufficient condition is that Theorem 2 provides a solution pair, given by (2), to the partition-dependent regulator equations (3). This pair is explicitly expressed in terms of the $k$-partition transformation term and the matrices satisfying the partition-independent regulator equations (3) and the LME (10). Thus, in the second design strategy to be given, each follower can recompute its feedforward gain $K_{2i}$ without recomputing a solution to any equations when the $k$-partition transformation term is updated.
Remark 10
Consider a MAS consisting of one leader and one follower. Let $\Phi$ (respectively, $\Theta$) be the set of leader and follower parameters that satisfy the conditions in Theorem 2 (respectively, (14)). Then, we have
$\Phi\setminus\Theta\neq\emptyset,\quad\Theta\setminus\Phi\neq\emptyset,\quad\Phi\cap\Theta\neq\emptyset.$ (15)
To see this, we first consider the following system parameters
$A_{0}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix},A_{1}=\begin{bmatrix}0&1\\ 0&0\end{bmatrix},B_{1}=\begin{bmatrix}0\\ 1\end{bmatrix},E_{1}=\begin{bmatrix}0&0\\ 0&0.5\end{bmatrix}$
[…]
IV-B Discussion: Synthesis of Control Parameters
This subsection first presents the second design strategy for the control parameters. Then, it compares both strategies from the perspectives of scalability and conservatism.
Remark 11
Let Conditions 2, 3, and 4 hold. Suppose that for any $i\in\mathcal{N}$, there is a solution to the LME (10). Per Theorem 2 and Corollary 1, the first two steps in Remark 6 remain unchanged, while the rest are updated as follows.
  1. (iii)
    For each $i\in\mathcal{N}$, find a pair $(X_{i},U_{i})$ that satisfies the partition-independent regulator equations (3).
  2. (iv)
    For each $i\in\mathcal{N}$, find a solution $Y_{i}$ to the LME (10).
  3. (v)
    For each $i\in\mathcal{N}$, let $K_{2i}=\tilde{U}_{i}-K_{1i}\tilde{X}_{i}$ where $\tilde{X}_{i}$ and $\tilde{U}_{i}$ are as defined in (2).
Remark 12
The second design strategy involves solving the partition-independent regulator equations (3) and the LME (10). Neither of these equations includes the $k$-partition transformation term $s_{i}$. Therefore, once each follower finds solutions to the partition-independent regulator equations (3) and the LME (10), it can use these solutions to recompute the feedforward gain $K_{2i}$ whenever $s_{i}$ is changed. As a result, the design strategy in Remark 11 is much more scalable than the one in Remark 6, as illustrated in Example 2.
Remark 13
Theorem 2 establishes that if the second design strategy is applicable, so is the first design strategy. Yet, the converse is not true. To see this, consider the MAS described in Remark 8 and let $\mathcal{F}=1$. Observe that the conditions in Remark 6 are satisfied. However, we conclude from Remark 8 that neither step (iii) nor step (iv) in Remark 11 is feasible. Consequently, the first design strategy applies to a broader class of leader and follower dynamics.
Lastly, Table I summarizes the differences between the first and second design strategies.
TABLE I: Differences in Design Strategies
Differences Strategies First Second     Conclusion
[…]
LMEs to be solved Partition-Dependent: (3) Partition-Independent:         (3) and (10)     The second is more scalable, as discussed in Remark 12.
IV-C Solvability of the Introduced LME
This subsection provides the straightforward characterizations of the solvability of the LME (10).
Proposition 1
Let $i\in\mathcal{N}$. Then the following conditions are equivalent:
  1. (i)
    There exists a solution $Y_{i}$ to the LME (10).
  2. (ii)
    The following inclusion holds:
    $\mathrm{im}\begin{bmatrix}E_{i}\\ G_{i}\end{bmatrix}\subseteq\mathrm{im}\begin{bmatrix}B_{i}\\ D_{i}\\ \end{bmatrix}.$ (16)
  3. (iii)
    The following rank condition holds:
    $\mathrm{rank}\begin{bmatrix}B_{i}&E_{i}\\ D_{i}&G_{i}\end{bmatrix}=\mathrm{rank}\begin{bmatrix}B_{i}\\ D_{i}\end{bmatrix}.$ (17)
Remark 14
The inclusion (16) is a structural characterization of the solvability of the LME (10). It is also easily testable by the rank condition (17). The inclusion (16) holds only when follower $i$ is subject to solely matched disturbances, which are common in certain applications (e.g., nonholonomic wheeled robots [33], UAVs [34], and spacecrafts [35]). Nevertheless, the first design strategy can still be employed in the presence of unmatched disturbances.
V Experimental and Numerical Illustrations
This section demonstrates the MORP’s flexibility in shifting mission objectives via an experiment with networked mobile robots. It also provides two numerical examples to showcase the MORP’s generality and compare the first and second design strategies regarding scalability. The following matrices are used throughout this section for the dynamics of the MASs.
$A_{\alpha}$ $=$ $\begin{bmatrix}0_{2\times 2}&I_{2}\\ 0_{2\times 2}&0_{2\times 2}\end{bmatrix},\ \Gamma=\begin{bmatrix}0&0.0025\\ -0.0025&0\end{bmatrix},$
$A_{\beta}$ $=$ $\begin{bmatrix}0.2&3\\ 0.1&-0.1\\ \end{bmatrix},\ B_{\beta}=\begin{bmatrix}0&3\\ 1&0\end{bmatrix}.$
Experiment 1
[…]
We assume that cylindrical and cuboid obstacles in the operating terrain are detected through a distributed sensor fusion algorithm that runs onboard each robot. Accordingly, the robots update their $k$-partition transformation terms. To simulate this, the $k$-partition transformation is defined as a piecewise constant function. Specifically, $S(t)=\mathrm{diag}(1,0.75,0.5)$ for $t\in[0,86.5)\cup[161,212]$ seconds and $S(t)=\mathrm{diag}(2.3,1.65,1)$ for $t\in[86.5,161)$ seconds. Note that Conditions 2, 3, 4, and the inclusion (16) hold. Following Remark 11, we set $\mu=10$ and design $K_{1i}$ using $place$ function in MATLAB [^5] For $i=1,2,3$, $\mathrm{spec}(A_{i}+B_{i}K_{1i})=\{-0.75,-1.25,-1.75,-2.5\}$.. As per steps (iii) and (iv), for $i=1,2,3$, a solution pair $(X_{i},U_{i})$ to the partition-independent regulator equations (3) is recovered from their equivalent system of linear equations (see the proof of Theorem 1.9 in [32]) and a solution $Y_{i}$ to the LME (10) is found using $linsolve$ function in MATLAB [^6] The authors thank Jackson Kulik for his comment on solving (10).. Lastly, based on $S$, we calculate $K_{2i}$ for each follower as in step (v).
The experiment is initiated with $x_{1}^{\mathrm{T}}(0)=[1.2,1.5,0,0]$, $x_{2}^{\mathrm{T}}(0)=[0.1,1.7,0,0]$, $x_{3}^{\mathrm{T}}(0)=[-0.5,1.3,0,0]$, $\eta_{i}(0)=0$ for $i=1,2,3$, and $v^{\mathrm{T}}(0)=[0,1]$. As seen in Fig. 1, the output of follower $i$ tracks the $s_{i}$ multiple of the leader’s state and steers around obstacles successfully for $i=1,2,3$. The experiment video can be found in the clickable link.
Figure 1: The trajectories of the MAS. Here, $y_{ij}$ denotes the $j$th entry of follower $i$’s output for $i=1,2,3$ and $y_{0j}$ denotes the $j$th entry of $v$ while “$\times$” and “o” marks $y_{ij}$ at the initial and final times for $i=0,1,2,3$, respectively.
Example 1
As indicated in Remark 2, the MORP includes the BORP. To make the MORP comparable with the existing solutions to the BORP, we force $k$-partition transformations to be gauge transformations. Despite such a restriction, this example presents the generality of the MORP. To this end, consider $100$ followers and a leader with the following matrices: $A_{i}=A_{\beta}$, $B_{i}=B_{\beta}$, $C_{i}=[1,0]$ for $i=1,\ldots,50$; $A_{i}=A_{\alpha}$, $B^{T}_{i}=[0\ I_{2}]$, $C_{i}=[1,0,0,0]$ for $i=51,\ldots,100$; $D_{i}=0$, $E_{i}=0$, $G_{i}=0$, $F_{i}=-1$ for $i=1,\ldots,100$; $A_{0}=0$. They communicate over $\mathcal{G}(\bar{\mathcal{A}})$ with $a_{i1}=1$ for $i=3,5,\ldots,99$, $a_{i1}=-1$ for $i=2,4,\ldots,100$, and $f_{1}=1$, while the remaining entries of $\mathcal{A}$ and $\mathcal{F}$ are zero.
As the bipartite framework partitions the followers [^7] There are studies incorporating the leader into the partition through the structural balance condition on $\mathcal{G}(\bar{\mathcal{A}})$ (e.g., see [15]). This, however, allows only $1$ BORP to be solved. based on the signs of the entries of $\mathcal{A}$, for the considered $\mathcal{G}({\mathcal{A}})$, it yields a unique $2$-partition of the followers. Therefore, the existing formulations in [17, 16, 18] allow only $2$ BORPs to be solved by swapping the followers that track the leader’s state and its additive inverse. On the other hand, as discussed in Section II-A, the number of $2$-partitions of the followers and $1$-partition of the followers obtained by $k$-partition transformations are respectively $12^{99}-$ 1 and $1$, which are the corresponding Stirling numbers of the second kind. In fact, there are $2^{100}$ gauge transformations generating all the aforementioned $2^{99}$ partitions. Thus, the proposed formulation allows $2^{100}$ BORPs to be solved without changing the underlying graph.
The simulation is initiated with $x_{i}^{\mathrm{T}}(0)=[3i/50,0]$ for $i=1,\ldots,50$, $x_{i}^{\mathrm{T}}(0)=[-3i/50+3,0,0,0]$, for $i=51,\ldots,100$, $\eta_{i}(0)=0$ for $i=1,\ldots,100$, and $v(0)=1$. The top row of Fig. 2 illustrates the output responses of 2 BORPs that can be solved using the bipartite framework and the proposed formulation, where $S=\pm\left(I_{50}\otimes\mathrm{diag}(1,-1)\right)$. For these BORPs, the design [^8] There is a typo in Equation (4b) of [17]. For Theorem 1 in [17] to be valid, the term $(z_{j}-\mathrm{sgn}(a_{ij})z_{i})$ needs to be replaced with $(\mathrm{sgn}(a_{ij})z_{j}-z_{i})$. in [17] and the first and second design strategies can generate identical output responses. The bottom row of Fig. 2 presents $2$ out of $22^{100}-$ 2 BORPs that can be solved with the proposed formulation but not with the bipartite framework under the same graph. Here, $S=\pm\left(\mathrm{diag}(1,-1)\otimes I_{50}\right)$.
Figure 2: The top row depicts the identical output responses with the design in [17] and the proposed design. The bottom row presents the output responses with the proposed formulation for 2 BORPs that are impossible to formulate with the approach in [17] without altering the underlying graph.
Example 2
This example compares the partition-dependent steps of both design strategies in terms of scalability. In particular, Fig. 3 shows the total elapsed times with an average laptop for follower 1 in Example 1 with $D_{1}=[0,1]$ and $E^{T}_{1}=[0,-4]$ to complete steps (iii) and (iv) of the first design (see Remark 6) and step (v) of the second design (see Remark 11) as the cardinality of the given set [^9] Such sets are generated using $randn$ function in MATLAB. of $k$-partition transformation terms increases. With the first design strategy, feedforward gains for up to $190$ $k$-partition transformation terms can be computed within $4$ milliseconds. On the other hand, with the second one, feedforward gains for approximately $10000$ $k$-partition transformation terms can be computed within the same amount of time.
Figure 3: Elapsed times of both design strategies with respect to the cardinality $M$ of the given set of $k$-partition transformation terms.
VI Conclusion
The primary motivation of this paper has been to provide MASs with objectives beyond cooperation and bipartition for tactical flexibility in adverse operating environments. To this end, the MORP for linear MASs has been formulated and solved for the first time. Two design strategies for the control parameters have been proposed. The first applies to a broader set of MASs, but it has a drawback due to the partition-dependent regulator equations. The second eliminates this drawback, and hence, it is significantly more scalable, yet applicable only when the followers are subject to matched disturbances. Table I summarizes the differences between the first and second design strategies. Theoretical results have been demonstrated by experimental and numerical tests.
The distributed observer in (III-A) assumes all followers have access to the matrix $A_{0}$. To relax this assumption to a small subset of the followers, solving the MORP with distributed control laws involving adaptive distributed observers (e.g., see [7]) would be well worth an exploration. To widen the application domain of the MORP, another research direction is investigating $k$-partition transformation generation as a distributed optimization problem, where the followers optimize a MAS-level objective such as total fuel or energy consumption.
Appendix
Lemma 3
Suppose Condition 4 holds. The matrix $A_{\mu}$ is Hurwitz if, and only if, $\mu$ satisfies the inequality (8).
Proof:
All the eigenvalues of $A_{\mu}$ are as follows:
$\lambda_{j}(A_{0})-\mu\lambda_{i}(\mathcal{H}),\quad j=1,\ldots,n_{0},\ i=1,\ldots,N$
(see the proof of Theorem 1 in [5]). One can use this fact and Lemma 1 in [5] to conclude the proof. ∎